OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 2, – 16), and x intercepts at x = -6 and x = 2. (Do not include the negative sign in your answer.)

OpenStudy (anonymous):

Do you want the general form or standard form?

OpenStudy (anonymous):

general hun

OpenStudy (anonymous):

(x + 6)(x - 2) = x^2 + 4x - 12 = 0 x^2 + 4x = 12 (x + 2)^2 = 16 y = (x + 2)^2 - 16